A butcher has some hamburger which is 4% fat and some hamburger which is 20% fat. How much 4% fat hamburger is needed to make 120 pounds of hamburger that is 10% fat?
Answer by htmentor(1343)

Let x = the number of pounds of 4% hamburger.
Then 120 - x will be the number of pounds of 20% hamburger
The number of pounds of fat in the 10% hamburger is 0.10*120.
So we need to solve the following equation for x:
0.04x + 0.20*(120 - x) = 0.10*120
0.04x + 24 - 0.2x = 12
Solving for x gives:
-0.16x = -12 -> x = 75
Ans: 75 pounds of 4% fat hamburger